OpenStudy (anonymous):

PLEASE HELP: Iron intake for a female adult is 3 mg more than twice the amount for a newborn. The amount for an adult female is 15 mg. How much should an infant receive?

OpenStudy (anonymous):

If the iron intake for a newborn should be x, and the iron intake of a female adult is 2x+3=15, then you can solve for x.

OpenStudy (anonymous):

Important that you read word for word. The problem says a female adult is (=) it never says a newborn is (=) so the newborn gets x newborn = x female adults = 2x + 3 The adult female = 15 so 2x + 3 = 15 Then you solve: And x will be the amount the newborn receives.

OpenStudy (anonymous):

<3 Thanks Everyone! My problem with these word problems is trying to figure out how to sort out the equation. Once i have it, its super easy to solve. If I get mixed up I stay stuck on it for hours! <3 i appreciate everyone's help! Thanks again!

OpenStudy (anonymous):

I find it helpful to define variables right away. If there is a subject of a sentence, it gets a variable letter. So in your Zebra problem, the Zebra's running speed gets a Z. The Cheetah's sprinting speed gets a C. And so on. Next I look at the relationships that are described between these variables. So the Cheetah's speed is three times the Zebra's speed minus 61. We have variables now, so this translates into a formula. C=3Z-61. Now I examine what is given, and substitute into my general formula. If the given is the Cheetah's speed, then I'm solving for Z. And so on. Practice, practice!